OpenStudy (anonymous):

Solve for the variable specified: V = 1wh: for w

OpenStudy (anonymous):

V = l*w*h (Initial formula) V = w (l*h) (Factor out to get w by itself) V/(l*h) = w (Divide both sides by l*h)

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If you're solving for variable w, what you're trying to accomplish is to isolate w by itself on one side of the equation, letting w equal some fixed number, which in this case is the volume divided by your box's length times width. This allows you to find the width of a box, in this case, knowing the length, height, and volume of the box, but not the width.